OpenStudy (butterflyhope):

*Will give Medal* Can you give me an example of an equation with a negative integer as the answer? Thanks!

OpenStudy (imstuck):

x - 3 = -10. When you solve for x, you get x = -7

OpenStudy (butterflyhope):

Thank you :)

OpenStudy (anonymous):

(supose x=3) 3-7 = 6-4 -4 = 2 x=2-4 x=-2